OpenStudy (anonymous):

Write each equation in Slope Intercept Form a) 2x-3y=9 b) 6x+9y=27

OpenStudy (chunkymonkay):

your in English lol

OpenStudy (anonymous):

for number one y=2/3x -3 and for second y=-2/3x - 3